SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Nov.29, 2005)

Attendees: Rainer Bartoldus, Kris Chaisanguanthum, Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su, Dan Walker.

Operations

  • We have been taking some physics data except yesterday and today (Nov.29) when PEP was busy making repairs. We took the opportunity to test a new version of the TSF firmware at IR2 yesterday. See more detailed in the DCT section below.

EMT

  • The EMT fastmon code monitoring the prescale factor will be in the next online base release. We cannot change the prescale factor to be greater than 1 until the tag is in the PR release. Dan will give Jim Hamilton a fastmon tag to install.

DCT/DCZ

  • The test_8 version of the TSF firmware passed all of the tests we could do at the test stand. Su Dong and Selina loaded this version at IR2 yesterday; however only the TSF-Y firmware was robust against all the tests. The new TSF-X firmware failed to output data to ZPD when the fast control cable was pulled. The old/original firmware also had this problem and corrupted DAQ data too when the fast control cable was pulled. We currently have the new firmware in the TSF-Y crate while the TSF-X crate still has the old firmware. Marc is re-routing the X version and keeping the Y version the same. We need to have test stand setup to include the interface to ZPD so that we can test for this failure mode.
  • Last Thursday (Nov.24), we got paged for a dip in SL10 around TSF-X15. See HN posted here. It lasted for a few runs in owl shift and into day shift. The dip disappeared after the detector was ramped down and then back up again. Chris O'Grady would like us to page him as well when this happened again. It was suggested that this is due to a problem with radiation reset. The DQM short-term instruction will be updated with this info.
  • Su Dong investigated further into the potential problem with the TSF inefficiencies seen in SL1,4,7,10. See HN posted here. One thing worths noting is that the charge distributions are very different for runs from April and November. It could have a significant impact on efficiency. There are some dips seen in the DCH hit efficiency vs wire for different superlayers. It's probably due to the new electronics. The DCH and ODF groups are aware of this, and will work to investigate this further. This might or might not explain the 0.5%-1% effect seen in the TSF efficiency vs superlayer plot.

IFT

  • One LST trigger board was pulled out from the detector on Monday to help investigate a problem with ramping up some of the LST channels. They found that the problem was still there without the trigger board in, so they inserted the board back in this morning. No physics runs were taken during the time the board was out, so no impact on anything at all.

GLT

  • Eric is generating some physics samples for L1 line trigger rate study.

Interface boards

  • Xuedong will help and teach Fang Fang and Selina how to set up the calibration framework to get it runs at the test stand as part of the firmware testing. Once we have it set up, we should be able to test 3 TSF boards with 1 ZPD board, and hopefully this will help us see the failure mode in the new TSF-X firmware. Documentations needed to set up for future reference.

L3

  • Rainer worked on re-structuring of the L3 DCH code. A new module named L3DTrackFinder was added. Eventually, we want to have track finding and fitting in different modules. The timing discrepancy was fixed now. It's in the nightly, and will be in release 18.6.3. Arik is working on the L3 tracking too.
  • Al suggested to add a plot of input rates vs time to the L3 monitoring JAS plots. There will be a trickle injection task force meeting about a week from now. It would be good to discuss and contribute to the trickle injection monitoring in that meeting.

Data Quality and Monitoring

  • Kris and Rahmat are working on strip charts, and they will show us some of them in our next meeting.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.